Grade 11Algebra

How to solve
4(m^3)-3m+1=0?
Please explain step wise.

Dishari Das
7 Years agoGrade 11

1 Answer

Shiva Pallikonda
7 Years ago
4m^3-3m+1=0
by verification.,m=[-1]satisfies the equation apply synthetic division method
-1| 4 0 -3 1
   |   -4  4  -1
   ---------------
     4 -4  1   0
   4m^2-4m+1=0
        [2m-1]^2=0
                   m=1/2
m=-1&1/2